# Quantum information with quantum-like bits

### v5

@@ VII Conclusions
In this paper, we built upon the general framework developed in Refs. Scholes, 2023, 2024; Scholes and Amati, 2025 to define and manipulate QL states. In ??, we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs. We introduced the Hilbert space for an arbitrary number of QL bits and explicitly derived the Bell states in two dimensions. In ??, we proposed a framework for defining quantum gates applicable to QL states and provided concrete examples of their action on many-body classical resources. ?? offers an overview of experimental platforms that could be used to implement and validate QL theory. Finally, in ??, we discussed that a QL notion of interference can emerge within our framework by presenting a set of axioms for interpreting QL states and observables.
@@
In this paper, we presented an early-stage analysis exploring the connection between QL gates and the many-body synchronization dynamics of oscillator systems. Specifically, we identified relationships between the structure of the adjacency matrix of the network and the synchronization patterns observed in the Kuramoto model. For future work, we aim to develop a more formal framework by deriving general expressions for multi-time correlation functions associated with network transformations induced by chains of QL gates. Additionally, we plan to investigate whether the structure of emergent QL states can be directly inferred from the synchronizing steady states and how this process is influenced by the microscopic details of the underlying classical dynamics. Future work could explore new possibilities regarding the generality of the present approach for different synchronizing dynamical systems. While this study focused on the nonlinear interactions of the Kuramoto model, it is crucial to assess the applicability of the results to a broader class of classical systems exhibiting synchronization. Extending this framework to more general models, where oscillators are represented by vectorial variables de Aguiar (2023) or elements of a unitary group, as in the Lohe model Lohe (2009); Scholes (2022), offers a promising path for further research. These generalizations enable the encoding of high-dimensional information within the phase relationships between graph nodes, potentially enhancing the modeling of complex dynamical systems.
In recent years, memcomputing has emerged as a new computational paradigm that leverages the intrinsic time-nonlocal properties of nonlinear classical dynamical systems to enable non-Boolean logic computations by exploiting memory effects within physical systems. A key feature of memcomputing is its terminal-agnostic nature, where input and output terminals are treated equivalently. This allows self-organizing circuits to dynamically adjust and converge to equilibrium points that correspond to solutions, regardless of the terminal through which data is introduced or extracted. Di Ventra and Traversa (2018) The feasibility of the memcomputing framework has been demonstrated across various physical implementations, including oscillator-based systems. Traversa et al. (2015); Di Ventra (2022) It can be insightful to implement a variant of our approach with flavors of memcomputing techniques, by exploiting time-nonlocality dissipative nonlinear dynamical systems, e.g. as predicted within the Nakajima–Mori–Zwanzig formalism. Nakajima (1958); Mori (1965); Zwanzig (1960); Dominy and Venturi (2017); Amati et al. (2022); Amati, Meyer, and Schilling (2019)
The environmental implications of large-scale industrial production of quantum processors remain a topic of ongoing debate. Ezratty (2023); Ikonen, Salmilehto, and Möttönen (2017); Paler and Basmadjian (2022); Sood and Chauhan (2024); Jaschke and Montangero (2023); Auffèves (2022) Conducting an early-stage assessment of the environmental impact of various resources can provide valuable metrics to identify the most energy-efficient and scalable quantum technologies for future development. In this context, implementing energy-efficient classical circuits for QL computing offers a promising pathway toward sustainable QL-processing engineering. While our perspective does not claim to provide a definitive solution, we recognize that classical systems, despite their larger size, could serve as viable energy-efficient alternatives for specific use cases, such as small-scale, proof-of-concept analyses in fundamental research.
@@ VIII Acknowledgments
We thank Debadrita Saha for useful discussions. This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
@@ Appendix A General transformation from emergent states to the computational basis
where $q=1,\cdots,{N_{\mathrm{QL}}}$, $i=1,\cdots,{N_{G}}$. ?? is expressed in matrix form as
@@
where $\left[D_{n}^{(q)}\right]_{ij}=\delta_{ij}a_{n,j}^{(q)}$, $n=1,2$. ?? can be inverted to determine a solution for $\left(x^{(q)},y^{(q)},w^{(q)},z^{(q)}\right)$ $\forall q=1,\cdots{N_{\mathrm{QL}}}$, corresponding to a unique expression for $U_{\mathrm{cb}}$.
@@ Appendix C Spectral properties of QL observables
    The linear operator $L(\phi)=\mathcal{F}\phi$, when restricted to $V_{1}$, is also Hermitian, and it admits an eigenvalue $\lambda_{2}$ with eigenvector $f_{2}\in V_{1}$. By construction, $f_{2}$ is orthogonal to $f_{1}$. We then define by $V_{2}$ the orthogonal complement to $\mathrm{span}\{f_{1},f_{2}\}$. Again, $\mathcal{F}$ maps $V_{2}$ onto itself. By proceeding this way, we find a sequence $\{\lambda_{n},f_{n},V_{n}\}_{n}$, where the subspace $V_{n}$ is orthogonal to $\mathrm{span}\{f_{1},\cdots,f_{n}\}$. The sequence ends at the step of order $2^{N_{\mathrm{QL}}}$, given that $\dim(V_{n})=2^{N_{\mathrm{QL}}}-n$. This process allows us to define a complete set of mutually orthogonal eigenvectors.
@@ Appendix D Dynamics of the Kuramoto model
or, more compactly as, $\dot{\bm{x}}=\tilde{\mathcal{R}}\bm{x}$, where $\bm{x}=\,\mathrm{e}^{i\bm{\theta}}=(\,\mathrm{e}^{i\theta_{1}},\cdots,\,\mathrm{e}^{i\theta_{2{N_{G}}}})^{\top}$ and ~$\tilde{\mathcal{R}}=\mathrm{diag}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}$. This allows us to obtain as a closed form for the solution of the dynamics of the Kuramoto model

### v6

@@ VII Conclusions
In this paper, we built upon the general framework developed in Refs. Scholes, 2023, 2024; Scholes and Amati, 2025 to define and manipulate QL states. In ??, we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs. We introduced the Hilbert space for an arbitrary number of QL bits and explicitly derived the Bell states in two dimensions. In ??, we proposed a framework for defining quantum gates applicable to QL states and provided concrete examples of their action on many-body classical resources. ?? offers an overview of experimental platforms that could be used to implement and test this theory. Finally, in ??, we discussed that a QL notion of interference can emerge within this framework by presenting a set of axioms for interpreting QL states and observables.
@@
As firstly discussed in ??, the number of edges in the Cartesian product graph defined in ??, along with the associated gate transformations in ??, increases exponentially with the number of QL bits. This exponential scaling imposes practical constraints on the system sizes that can be investigated with current computational and experimental capabilities. Nevertheless, we believe that moderately sized networks are sufficient to demonstrate the core principles of QL algorithms. For example, entangled states such as Bell and GHZ states can be effectively generated and analyzed using shallow-depth circuits. Scholes and Amati (2025) In terms of scaling, to date high-performance classical simulations allow the implementation of systems with up to approximately ${N_{\mathrm{QL}}}\lesssim 5-10$ QL bits and ${N_{G}}\lesssim 10$. Simulations of a Kuramoto model with this system size are given, e.g., in Ref. Skardal, 2018.
@@
The scope and applicability of our approach could be broadened by extending the treatment of quantum gates to any unitary transformation, in particular to time evolution operators. From this perspective, one direction for future research is to map the dynamics of a unitary quantum subsystem onto a large network of classical oscillators and, for instance, to explore the effects of decoherence arising from interactions with an external dissipative environment. Accurately reproducing quantum time evolution through classical dynamics remains a nontrivial challenge. It requires to extend our approach such that of the algebra of Hilbert space generators is preserved, at least approximately. Runeson and Richardson (2019); Kapral (2015) This strategy becomes particularly promising when coupled with the enforcement of quantum detailed balance at long times—a direction the authors have recently explored in the context of nonadiabatic quantum dynamics. Amati, Mannouch, and Richardson (2023); Amati, Runeson, and Richardson (2023)
The extent to which our framework offers any advantage compared to quantum computing remains an open question, that we believe will be fully explored by testing further approximations of the exact mapping proposed here. More specifically, our approach requires an exponentially large number of oscillators to encode the correlations of the Hilbert space of a system of QL bits. This high resource demand does not offer per se any advantage in regards to scaling. In Appendix A of our new paper, Ref. Amati and Scholes, 2025, we outline as a potential strategy to mitigate the high cost of our approach by replacing full graph representations with their lower rank approximations. This can lead to a renormalization of both graph resources and corresponding gates, possible by truncating negligible parts of the spectrum. In general, we believe that it can be worth challenging the trade-off between implementing genuine quantum resources with polynomial complexity and easier accessible classical resources with unfavorable exponential scaling. We view these observations as a starting point for further investigations into the foundational limits and capabilities of our approach and quantum-inspired frameworks in general.
Future work could explore new possibilities regarding the choice of resources used for information processing. For example, our framework could be extended to other flavors of Kuramoto-type models, where oscillators are represented by vectorial variables de Aguiar (2023) or elements of a unitary group. Lohe (2009); Scholes (2022) Other classical systems could be used as a resource for information processing in a similar fashion as what we proposed here. An example in this regard is the Fermi–Pasta–Ulam–Tsingau chain, where isolated solitary waves, emerging from symplectic classical nonlinear dynamics, could be used to store and process information in a QL fashion. Amati, Meyer, and Schilling (2019); Carati et al. (2015); Amati and Schilling (2020); Berman and Izrailev (2005) Since the dynamics of this classical model are Hamiltonian, the approach may allow for a well-defined classical detailed balance condition, which could be enforced to match its quantum analog and ensure correct long-time thermalization. Amati, Mannouch, and Richardson (2023)
@@ VIII Acknowledgments
The authors thank Dr. Debadrita Saha for useful discussions. This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
@@ Appendix A General transformation from emergent states to the computational basis
where $q=1,\ldots,{N_{\mathrm{QL}}}$, $i=1,\ldots,{N_{G}}$. ?? is expressed in matrix form as
@@
where $\left[D_{n}^{(q)}\right]_{ij}=\delta_{ij}a_{n,j}^{(q)}$, $n=1,2$. ?? can be inverted to determine a solution for $\left(x^{(q)},y^{(q)},w^{(q)},z^{(q)}\right)$ $\forall q=1,\ldots{N_{\mathrm{QL}}}$, corresponding to a unique expression for $U_{\mathrm{cb}}$.
@@ Appendix C Spectral properties of QL observables
    The linear operator $L(\phi)=\mathcal{F}\phi$, when restricted to $V_{1}$, is also Hermitian, and it admits an eigenvalue $\lambda_{2}$ with eigenvector $f_{2}\in V_{1}$. By construction, $f_{2}$ is orthogonal to $f_{1}$. We then define by $V_{2}$ the orthogonal complement to $\mathrm{span}\{f_{1},f_{2}\}$. Again, $\mathcal{F}$ maps $V_{2}$ onto itself. By proceeding this way, we find a sequence $\{\lambda_{n},f_{n},V_{n}\}_{n}$, where the subspace $V_{n}$ is orthogonal to $\mathrm{span}\{f_{1},\ldots,f_{n}\}$. The sequence ends at the step of order $2^{N_{\mathrm{QL}}}$, given that $\dim(V_{n})=2^{N_{\mathrm{QL}}}-n$. This process allows us to define a complete set of mutually orthogonal eigenvectors.
@@ Appendix D Dynamics of the Kuramoto model
or, more compactly as, $\dot{\bm{x}}=\tilde{\mathcal{R}}\bm{x}$, where $\bm{x}=\,\mathrm{e}^{i\bm{\theta}}=(\,\mathrm{e}^{i\theta_{1}},\ldots,\,\mathrm{e}^{i\theta_{2{N_{G}}}})^{\top}$ and ~$\tilde{\mathcal{R}}=\mathrm{diag}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}$. This allows us to obtain as a closed form for the solution of the dynamics of the Kuramoto model